# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v2

@@ IV-A Composite detection and response matrix estimation
The main objective of typical radar systems is to detect targets and estimate the matrix response for each one. After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
$\{\hat{\mathbf{H}}_{t},\hat{\mathcal{H}}_{q}\}=\arg\max_{\mathbf{H}_{t},\mathcal{H}_{q}\forall q\in\{0,1\}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{q}).$ (27)
This joint optimisation problem aims to simultaneously detect the presence of a target (determining ^$\hat{\mathcal{H}}_{q}$) and estimate its response matrix ($\hat{\mathbf{H}}_{t}$) if present. For the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, the PDF of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis can be shown as a multivariate normal density,
$f(\mathbf{y}_{\mathrm{r},t,l};\,$ $\mathbf{H}_{t}q,\,\mathbf{x}_{t}q,\,\mathcal{H}_{q})=\frac{1}{\pi^{N}\left[(1-q)\sigma_{n}^{2N}+q\,\det(\mathbf{R}_{1,t})\right]}$
@@
where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$. The generalised likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses, derived as,
$\Lambda_{t,l}=\frac{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{1})}{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathcal{H}_{0})}.$ (29)
As the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$, the target response matrix is estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem we obtain,
$\hat{\mathbf{H}}_{t}=\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t},\mathcal{H}_{1}).$ (30)
The likelihood function is Gaussian, and the maximisation leads to the least squares solution,
$\hat{\mathbf{H}}_{t}$ $=\arg\min_{\mathbf{H}_{t}}\left\|\mathbf{y}_{\mathrm{r},t}-\mathbf{H}_{t}\mathbf{x}_{t}\right\|^{2}=\mathbf{y}_{\mathrm{r},t}\mathbf{x}_{t}^{H}\left(\mathbf{x}_{t}\mathbf{x}_{t}^{H}\right)^{-1}.$ (31)
This least squares solution provides an unbiased estimate of $\mathbf{H}_{t}$. Substituting the PDF in (28) into the GLRT, we obtain,
@@
$\frac{\pi^{N}\sigma_{n}^{2N}\exp\left(-(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})\right)}{\pi^{N}\det(\hat{\mathbf{R}}_{1,t})\exp\left(-\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}\right)}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{0}$ (32)
where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold. Taking the logarithm and rearranging, we obtain,
$\ln\Lambda(\mathbf{y}_{\mathrm{r},t,l})$ $=\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}-N\ln\sigma_{n}^{2}+\ln\det(\hat{\mathbf{R}}_{1,t})$
$\!-(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}-\hat{\mathbf{H}}_{t}\mathbf{x}_{t})\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\ln\tau_{0}$ (33)
The terms $N\ln\sigma_{n}^{2}$ and $\ln\det(\hat{\mathbf{R}}_{1,t})$ are constant with respect to $\mathbf{y}_{\mathrm{r},t,l}$, so they can be absorbed into the threshold, which further simplifies our GLRT as follows,
@@
where $\tau_{1}=\ln\tau_{0}+N\ln\sigma_{n}^{2}-\ln\det(\hat{\mathbf{R}}_{1,t})$. This expression compares hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$. To use all $L$ snapshots with independent signals, we average as,
$\bar{\Lambda}_{t,l}=\frac{1}{L}\sum_{l=1}^{L}\dot{\Lambda}_{t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{2}.$ (35)
where $\tau_{2}$ is the detection threshold considering all snapshots and can be designed per the Neyman-Pearson lemma.
@@ IV-B The KLD of radar system
where $\mathbf{R}_{2,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{1,t}=\mathbf{H}_{t}\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}=\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathrm{det}(.)$ is the determinant operation. $\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$.
@@ V Radar Waveform Optimisation
s.t. $\underline{A}_{t}\leq{\mathrm{{KLD}}_{\mathrm{r},t}},\,\,\,\forall t\in T,$ (37b)
$\underline{B}_{k}\leq{\mathrm{{KLD}}_{\mathrm{c},k}},\forall k\in K,$ (37c)
@@
where $\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ is defined in (36). After substituting (36) in (37a), can be written as $\mathcal{P}_{2}$ that is given on page V. This optimisation problem focuses on the radar waveform while considering the whole ISAC system performance. The constraint on radar power (37d) is explicit because the radar waveform is our primary optimisation variable. The communication power is implicitly considered through the ZF beamforming scheme and the constraint (37c) on communication performance. Specifically, the communication subsystem utilises its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information.
@@
$+\mathrm{Tr}\left(\left(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}\right)^{-1}\>\left(\sigma_{n}^{2}\>\mathbf{I}_{N}\right)\right)-N\>\left(1+\ln\left(\sigma_{n}^{2}\right)\right)\Bigg{)}$
@@
where $\mathbb{I}_{1}=\mathbb{I}(\mathrm{KLD}_{\mathrm{r},t}<\underline{A}_{t})$ and $\mathbb{I}_{2}=\mathbb{I}(\mathrm{KLD}_{\mathrm{c},k}<\underline{B}_{k})$ with $\mathbb{I}(\cdot)$ represents the indicator function. The gradients of $\mathrm{KLD_{r}}$ and $\mathrm{KLD_{c}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix I.A. The final forms are given by,
@@
$[\nabla\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}},\mathbf{H}_{t}}]_{i,l,t}=$ $-\frac{2\>d_{\mathrm{c},k}^{\zeta}\lambda p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.$ (44)
@@
where $\gamma>1$ is a constant factor. This adaptive scheme increases the penalty when constraints are violated and keeps it constant when they are satisfied.
@@
The projection operator $\mathcal{P}$ ensures that the power constraint is always satisfied. The backtracking line search procedure is used to determine an appropriate step size $\alpha_{n}$ at each iteration. This adaptive step size selection helps to ensure a sufficient increase in the objective function while maintaining the stability of the algorithm. The line search condition can be expressed as,
$\begin{split}\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)})\leq&\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)})\\ &+c\alpha_{n}\langle\nabla\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}),\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}-\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\rangle,\end{split}$ (49)
@@ VI Communication optimisation
The communication subsystem optimisation uses the gradient-assisted IPM [39], in contrast to the projected gradient method with penalty used for radar optimisation, due to the different optimisation variable structures. Radar optimisation involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, while communication beamforming optimises a two-dimensional matrix $\mathbf{W}_{c}\in\mathbb{C}^{N\times K}$. This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions. The communication waveform is optimised while using the conventional identity covariance matrix design for the radar subsystem, with $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$. The optimisation problem can be expressed as follows,
$\mathcal{P}_{3}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (50a)
s.t. $\underline{A}_{t}\leq{\mathrm{{KLD}}_{\mathrm{r},t}},\,\,\,\forall t\in T,$ (50b)
$\underline{B}_{k}\leq{\mathrm{{KLD}}_{\mathrm{c},k}},\forall k\in K,$ (50c)
@@
$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}\quad\text{for all }j$ (52)
@@
$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.$ (55)
@@
$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.$ (56)
@@
$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).$ (57)
@@
0:  Initial point $\mathbf{W}_{c}^{(0)}$, constants $\mu_{0}$, $\gamma$, $P_{\mathrm{c}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\epsilon$, and maximum iterations $\mathrm{max\_iter}$
0:  Optimal solution $\mathbf{W}_{c}^{*}$
@@
3:     Compute $f(\mathbf{W}_{c}^{(n)})$, $\nabla f(\mathbf{W}_{c}^{(n)})$, $g_{i}(\mathbf{W}_{c}^{(n)})$, and $\nabla g_{i}(\mathbf{W}_{c}^{(n)})$
@@
8:     Solve Newton system for search direction $\Delta\mathbf{W}_{c}$
@@
10:     Update $\mathbf{W}_{c}^{(n+1)}=\mathbf{W}_{c}^{(n)}+\alpha\Delta\mathbf{W}_{c}$
@@
14:  return  $\mathbf{W}_{c}^{*}=\mathbf{W}_{c}^{(n)}$

### v4

@@ IV-A Composite detection and response matrix estimation
The main objective of typical radar systems is to detect the existence of targets and estimate the matrix response for existing targets. After separating signals associated with different targets, we formulate a composite detection-estimation problem as follows,
$\{\hat{\mathbf{H}}_{t},\hat{\mathcal{H}}_{q}\}=\arg\max_{\mathbf{H}_{t},\mathcal{H}_{q}\forall q\in\{0,1\}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{q}).$ (19)
This joint optimisation problem aims to simultaneously detect the presence of a target, i.e., determining ^$\hat{\mathcal{H}}_{q}$, and estimate its response matrix $\hat{\mathbf{H}}_{t}$ if presents. It can be observed from (5) that the distribution of the received signals vector $\mathbf{y}_{\mathrm{r},t,l}$ under each hypothesis is multivariate normal whose PDF is,
$f(\mathbf{y}_{\mathrm{r},t,l};\,$ $\mathbf{H}_{t}q,\,\mathbf{x}_{t,l}q,\,\mathcal{H}_{q})=\frac{1}{\pi^{N}\left[(1-q)\sigma_{n}^{2N}+q\,\det(\mathbf{R}_{1,t})\right]}$
@@
where $\mathbf{R}_{1,t}=\mathbf{H}_{t}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathbf{R}_{t,l}=\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}\>d_{\mathrm{r},t}^{\zeta}$. The generalised likelihood ratio test (GLRT) for the $t$-th target is used to determine target existence or absence by comparing the likelihoods under the two hypotheses, which is derived as,
$\Lambda_{t,l}=\frac{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{1})}{\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathcal{H}_{0})}.$ (21)
As the PDF does not depend on $\mathbf{H}_{t}$ under $\mathcal{H}_{0}$, the target response matrix is estimated under $\mathcal{H}_{1}$, thus by using the maximum likelihood estimation theorem, we obtain,
$\hat{\mathbf{H}}_{t}=\arg\max_{\mathbf{H}_{t}}f(\mathbf{y}_{\mathrm{r},t,l};\mathbf{H}_{t},\mathbf{x}_{t,l},\mathcal{H}_{1}).$ (22)
Since the PDF of $\mathbf{y}_{\mathrm{r},t,l}$ is multivariate Gaussian, the maximisation leads to the least squares solution given by,
$\hat{\mathbf{H}}_{t}\!$ $=\!\arg\min_{\mathbf{H}_{t}}\left\|\mathbf{y}_{\mathrm{r},t}\!-\!\mathbf{H}_{t}\mathbf{x}_{t,l}\right\|^{2}\!=\!\mathbf{y}_{\mathrm{r},t}\mathbf{x}_{t,l}^{H}\left(\mathbf{x}_{t,l}\mathbf{x}_{t,l}^{H}\right)^{-1}.$ (23)
This least squares solution provides an unbiased estimate of the target‐response matrix $\mathbf{H}_{t}$, with accuracy directly tied to the detection probability $P_{D}$. When the radar correctly detects a target under $\mathcal{H}_{1}$, the maximum likelihood solution $\hat{\mathbf{H}}_{t}$ is used to estimate $\mathbf{H}_{t}$. In contrast, a miss detection (i.e., deciding $\mathcal{H}_{0}$ under $\mathcal{H}_{1}$) forces $\hat{\mathbf{H}}_{t}=\mathbf{0}$, resulting in significant errors. The overall unconditional MSE is defined as,
$\mathrm{MSE}$ $=p(H_{1})\Bigl{[}(1-P_{D})\,\mathrm{MSE}_{0\_1}+P_{D}\,\mathrm{MSE}_{1\_1}\Bigr{]}$
$+\bigl{[}1-p(H_{1})\bigr{]}\Bigl{[}P_{\mathrm{FA}}\,\mathrm{MSE}_{1\_0}\Bigr{]},$ (24)
[…]
Substituting the PDF in (20) into the GLRT, we obtain,
@@
$\frac{\!\sigma_{n}^{2N}\!\exp\!\left(\!-(\mathbf{y}_{\mathrm{r},t,l}\!-\!\hat{\mathbf{H}}_{t}\mathbf{x}_{t,l})^{H}\hat{\mathbf{R}}_{1,t}^{-1}(\mathbf{y}_{\mathrm{r},t,l}\!-\!\hat{\mathbf{H}}_{t}\mathbf{x}_{t,l})\!\right)}{\!\det(\hat{\mathbf{R}}_{1,t})\exp\!\left(\!-\frac{1}{\sigma_{n}^{2}}\mathbf{y}_{\mathrm{r},t,l}^{H}\>\mathbf{y}_{\mathrm{r},t,l}\!\right)}\!\!\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\!\tau_{0}$ (25)
where $\hat{\mathbf{R}}_{1,t}=\hat{\mathbf{H}}_{t}\mathbf{R}_{t,l}\hat{\mathbf{H}}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$ and $\tau_{0}$ is the detection threshold. Taking the logarithm of the likelihood ratio and rearranging terms, the final GLRT form is given by,
@@
where $\tau_{1}=\ln\tau_{0}+N\ln\sigma_{n}^{2}-\ln\det(\hat{\mathbf{R}}_{1,t})$. Here, the constant terms $N\ln\sigma_{n}^{2}$ and $\ln\det(\hat{\mathbf{R}}_{1,t})$ have been absorbed into the threshold. This expression decides the most probable hypothesis, i.e., $\mathcal{H}_{0}$ or $\mathcal{H}_{1}$, by incorporating the estimated target response matrix $\hat{\mathbf{H}}_{t}$ in the covariance matrix $\mathbf{R}_{1,t}$. Since the received signals from different snapshots are conditionally independent, (26) can be updated to adapt to $L$ snapshot scenario by taking the average, i.e., $\bar{\Lambda}_{t,l}=\frac{1}{L}\sum_{l=1}^{L}\dot{\Lambda}_{t,l}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\tau_{2}$, with $\tau_{2}$ is the detection threshold considering all snapshots and can be designed according to Neyman-Pearson lemma.
@@ IV-B The KLD of radar system
where $\mathbf{R}_{2,t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{1,t}=\mathbf{H}_{t}\frac{1}{L}\sum_{l=1}^{L}\mathbf{R}_{t,l}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}=\mathbf{H}_{t}\mathbf{R}_{t}\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}$, and $\mathrm{det}(.)$ is the determinant operation. $\mathrm{KLD}_{\mathrm{r},t}=\frac{1}{2}\left(\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}+\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}\right)$, where in this case $\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{0}\rightarrow\mathcal{H}_{1}}=\mathrm{KLD}_{\mathrm{r},t}^{\mathcal{H}_{1}\rightarrow\mathcal{H}_{0}}$. The average KLD of the whole ISAC system can be formulated as follows,
$\mathrm{KLD}_{\mathrm{ISAC}}=\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k}+\frac{1}{T}\sum_{t=1}^{T}\mathrm{KLD}_{\mathrm{r},t}.$ (28)
@@ V Radar Waveform Optimisation
s.t. ${\mathrm{{KLD}}_{\mathrm{r},t}}\geq\underline{A}_{t},\,\,\,\forall t\in T,$ (29b)
${\mathrm{{KLD}}_{\mathrm{c},k}}\geq\underline{B}_{k},\forall k\in K,$ (29c)
@@
where $\mathrm{KLD}_{\mathrm{r},t|\mathbf{W}_{\mathrm{r},t},\mathbf{H}_{t}}$ is defined in (27). After substituting (27) in (29a), can be written as $\mathcal{P}_{2}$ that is given on page V. This optimisation problem focuses on the radar waveform while considering the whole ISAC system performance. The minimum KLD requirements in constraints (29b) and (29c) are carefully chosen to ensure specific performance guarantees in terms of detection probability $P_{D}$ and BER for the radar and communication subsystems, respectively. The constraint on radar power (29d) is explicit because the radar waveform is our primary optimisation variable. The communication power is implicitly considered through the ZF beamforming scheme and the constraint (29c) on communication performance. Specifically, the communication subsystem utilises its allocated power through the ZF beamforming scheme to distribute power to each UE, with weights predetermined based on the channel state information.
@@
$+\mathrm{Tr}\left(\left(\>d_{\mathrm{r},t}^{\zeta}\mathbf{H}_{t}(\frac{1}{L}\mathbf{W}_{\mathrm{r},t}\mathbf{W}_{\mathrm{r},t}^{H}+\mathbf{W}_{\mathrm{c}}\mathbf{W}_{\mathrm{c}}^{H})\mathbf{H}_{t}^{H}+\sigma_{n}^{2}\mathbf{I}_{N}\right)^{-1}\!\sigma_{n}^{2}\!\right)\!-\!N\!\left(1+\ln\left(\sigma_{n}^{2}\right)\right)\!\Bigg{)}$
@@
where $\mathbb{I}_{1}=\mathbb{I}(\mathrm{KLD}_{\mathrm{r},t}<\underline{A}_{t})$ and $\mathbb{I}_{2}=\mathbb{I}(\mathrm{KLD}_{\mathrm{c},k}<\underline{B}_{k})$ with $\mathbb{I}(\cdot)$ represents the indicator function. The gradients of $\mathrm{KLD_{\mathrm{r}}}$ and $\mathrm{KLD_{\mathrm{c}}}$ with respect to $\mathbf{W}_{\mathrm{r},t}$ and $\mathbf{\dot{W}}_{\mathrm{r}}$, respectively, are provided in Appendix I.A. The final forms are given by,
@@
$[\nabla\mathrm{KLD}_{\mathrm{c},k|\mathbf{\dot{W}}_{\mathrm{r}},\mathbf{H}_{t}}]_{i,l,t}\!=\!$ $-\!\frac{2d_{\mathrm{c},k}^{\zeta}\lambda p_{k}(N\!\!-\!\!K)\sigma_{h}^{2}}{M(M\!\!-\!\!1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\!\mathrm{r}}]_{i,l,t}.$ (36)
@@
where $\gamma>1$ is a constant factor. This adaptive scheme increases the penalty when constraints are violated and keeps it constant when they are satisfied. The projection operator $\mathcal{P}$ ensures that the power constraint is always satisfied. The backtracking line search procedure is used to determine an appropriate step size $\alpha_{n}$ at each iteration. This adaptive step size selection helps to ensure a sufficient increase in the objective function while maintaining the stability of the algorithm. The line search condition can be expressed as,
$\begin{split}\!\!\phi(\mathbf{\dot{W}}_{\mathrm{\!r}}^{(n+1)})\!\!\leq&\phi(\mathbf{\dot{W}}_{\mathrm{\!r}}^{\!(n)})\!+\!c\alpha_{n}\!\langle\nabla\!\phi(\mathbf{\dot{W}}_{\mathrm{r}}^{\!(n)}),\!\mathbf{\dot{W}}_{\mathrm{r}}^{(n+1)}\!\!-\!\!\mathbf{\dot{W}}_{\mathrm{r}}^{(n)}\rangle,\!\!\!\end{split}$ (41)
@@ VI Communication optimisation
The communication subsystem optimisation uses the gradient-assisted IPM [43], in contrast to the projected gradient method with penalty used for radar optimisation, due to the different optimisation variable structures. Radar optimisation involves a three-dimensional tensor $\mathbf{\dot{W}}_{\mathrm{r}}\in\mathbb{C}^{N\times L\times T}$, while communication beamforming optimises a two-dimensional matrix $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$. This reduced dimensionality makes IPM particularly suitable, potentially offering faster convergence and more precise solutions. The communication waveform is optimised while using the conventional identity covariance (CIC) matrix design for the radar subsystem, with $\mathbf{R}_{w}=\mathbf{I}_{N\times N}\quad\forall t\in T$. The optimisation problem can be expressed as follows,
$\mathcal{P}_{4}:\max_{\mathbf{W}_{\mathrm{c}}}\quad\!\!$ $\frac{1}{K}\sum_{k=1}^{K}\mathrm{KLD}_{\mathrm{c},k|\mathbf{W}_{\mathrm{c}},\mathbf{h}_{k}}$ (42a)
s.t. ${\mathrm{{KLD}}_{\mathrm{r},t}}\geq\underline{A}_{t},\,\,\,\forall t\in T,$ (42b)
${\mathrm{{KLD}}_{\mathrm{c},k}}\geq\underline{B}_{k},\forall k\in K,$ (42c)
@@
$\nabla_{\mathbf{w}_{\mathrm{c},j}}|\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j}|^{2}$ $=2d_{\mathrm{c},k}^{\zeta}\mathbf{h}_{k}\mathbf{h}_{k}^{T}\mathbf{w}_{\mathrm{c},j},\quad\quad\quad\quad\forall j$ (44)
@@
$g(\mathbf{W}_{\mathrm{c}})=\|\mathbf{W}_{\mathrm{c}}\|_{F}^{2}-P_{\mathrm{c}}\leq 0.\vspace{-0.05 in}$ (47)
@@
$\nabla_{\mathbf{w}_{\mathrm{c},j}}g(\mathbf{W}_{\mathrm{c}})=2\mathbf{w}_{\mathrm{c},j}.\vspace{-0.05 in}$ (48)
@@
$\nabla_{\mathbf{W}_{\mathrm{c}}}\mathrm{KLD}_{\mathrm{r},t}=\frac{d_{\mathrm{r},t}^{\zeta}}{\ln 2}\left(\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{c}}\mathbf{H}_{t}^{H}\right).\vspace{-0.03 in}$ (49)
@@
0:  Initial point $\mathbf{W}_{\mathrm{c}}^{(0)}$, constants $\mu_{0}$, $\gamma$, $P_{\mathrm{c}}$, $\underline{A}_{t}$, $\underline{B}_{k}$, tolerance $\epsilon$, and maximum iterations $\mathrm{max\_iter}$
0:  Optimal solution $\mathbf{W}_{\mathrm{c}}^{*}$
@@
3:     Compute $f(\mathbf{W}_{\mathrm{c}}^{(n)})$, $\nabla f(\mathbf{W}_{\mathrm{c}}^{(n)})$, $g_{i}(\mathbf{W}_{\mathrm{c}}^{(n)})$, and $\nabla g_{i}(\mathbf{W}_{\mathrm{c}}^{(n)})$
@@
8:     Solve Newton system for search direction $\Delta\mathbf{W}_{\mathrm{c}}$
@@
10:     Update $\mathbf{W}_{\mathrm{c}}^{(n+1)}=\mathbf{W}_{\mathrm{c}}^{(n)}+\alpha\Delta\mathbf{W}_{\mathrm{c}}$
@@
14:  return  $\mathbf{W}_{\mathrm{c}}^{*}=\mathbf{W}_{\mathrm{c}}^{(n)}$